OpenStudy (jiteshmeghwal9):

Formula of nitrate of a metal is \(\Large{M(NO_3)_3}\); the formula of its pyrophosphate is :- A. \(\Large{M_3(P_2O_7)_4}\) B. \(\Large{M(P_2O_7)_2}\) C. \(\Large{M_2(P_2O_7)_3}\) D. \(\Large{M_4(P_2O_7)_3}\)

OpenStudy (jiteshmeghwal9):

@Chlorophyll @chemistry_help plz help:)

OpenStudy (australopithecus):

Here is a picture of pyrophosphate

OpenStudy (australopithecus):

First determine what the charge of NO3 is

OpenStudy (anonymous):

@Australopithecus can u plz help me with my chem question?

OpenStudy (australopithecus):

its -1, so you have -3 thus M has to have a charge of +3 to render the molecule neutral

OpenStudy (jiteshmeghwal9):

metals always loses electrons therefore it must have a positive charge.

OpenStudy (australopithecus):

notice that Pyrophosphate has a charge of -4

OpenStudy (anonymous):

Formula: P2O7^4−

OpenStudy (australopithecus):

What ratio of M and pyrophosphate do you need to render a neutral molecule

OpenStudy (australopithecus):

HINT: Remember 4*3 = 12

OpenStudy (australopithecus):

do you follow?

OpenStudy (jiteshmeghwal9):

sorry i m confusing :(

OpenStudy (jiteshmeghwal9):

confused*

OpenStudy (australopithecus):

NO3 has a charge of -1 (NO3)3 means that there are three groups of this polyatomic ion thus there is a total charge given by them of -3 so you have (NO3)3 -3 thus M has to have a charge of +3 because the molecule is neutral. We now know the charge of M is +3 We also know that pyrophosphate has a charge of -4 thus we need to figure out the ratio of M to Pyrophosphate to render a neutral compound

OpenStudy (australopithecus):

I gave you the hint 3*4 = 12 think about it, you need to have an amount of M that gives a charge equal to the amount of pyrophosphate

OpenStudy (anonymous):

@jiteshmeghwal9 Do you agree up the charge of M is +3?

OpenStudy (jiteshmeghwal9):

yes

OpenStudy (anonymous):

Since from the negative charge, we have P2O7 charge is -4!

OpenStudy (anonymous):

-> By the rule, total charge must be neutral, how do we balance between -4 and +3, so they will cancel out?

OpenStudy (jiteshmeghwal9):

I have not studied Chemistry other than Atomic Structure so sorry i can't do this man. I'm going

OpenStudy (australopithecus):

+3 * 4 = 12 -4 * 3 = 12 so we need 4 molecules of M and we need 3 molecules of Pyrophosphate thus we have M4(P2O7)3

OpenStudy (australopithecus):

+3 * 4 = +12 charge -4 * 3 = -12 charge